Shift summary for 20 Sept. 1999 owl

Ashkan_Alavi-Harati (alavi@fnal.gov)
Mon, 20 Sep 1999 07:10:46 -0500

Shift Summary for owl shift, 20/Sept/99

On Shift: Ashkan, Valeri

Hours of beam available at all intensities: 4.0
Hours of beam available at nominal intensity: 1.0
Hours of E799 data (special runs not included): 2.8
Hours of special runs (laser scans, muon runs...): 0.2

Accounting of lost time: Accelerator various problems, intermittent low
intensity beam due to F17 kicker problem. Kumquat problems,
run getting in weird state.

========================================================================
Runs ended this shift:
14635 BEAM_799 4.8M events.
14640 BEAM_MU_4E12 0.35M events.
14642 BEAM_799 1.5M events.
14643 BEAM_799 ongoing.

Tapes brought to FCC this shift: none.

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: smooth E799 data taking for the first hour, low
intensity intermittent beam or no beam for the rest of the shift.
A muon run also taken.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14635 299 00:16 ktevana errors, KFREND.
301 00:18 ktevana errors, KFREND.
330 01:00 Stop the run. No beam for the last two spills.
Call MCR to set the magnets for muons.
Update ADGN14635 and DCTO14635 in the DB.
01:03 F17 kicker problems, intermittent beam to users.
01:05 Beam gone.
01:20 Initialization for the muon run hanging at stage
3, complaining about pipeline crate 6 and kumquats.
dart_fresh_start didn't resolve the kumquat error.
Reloading kumquat maps hanged ktrig with
tcl error messages.
01:35 Quench at A4, F2 and A7, estimate 45 min.
02:00 Valeri is taking advantage of no-beam situation to
solve the kumquat trigger problem before waking
Jason up. Manual fscc resetting followed by
reloading maps and dart_fresh_start didn't work
again. Finally page Jason.
02:14 Low intensity beam (3.2E12) back , the kicker
problem still exists, beam will be intermittent.

02:25 Ashkan made new XT maps from run 14635 and
installed in the database.

02:30 Jason gets to the hall to work on the kumquats.
The problem seemed to be fixed by taking the gate
off and putting it back followed by a power cycle.

14640 2 02:50 Finally the initialization goes through, start a
muon run.
7 03:00 Stop the run. This is the first muon run since the
plateau studies at beginning of E799. Flipped the
analysis magnet to Neg polarity. Ask for Hadrons.
The intensity is now above 7E12 but still
intermittent beam.

03:10 Trigger initialization fails again at kumquat
stage. Jason back to work!

03:36 Yet another quench, estimate 30 min.

14642 1 03:50 Jason fixes the problem by replacing a module.
Start the new beam_799 run.

20 04:15 Beam is back, SEM=7.4E12

?? 05:20 Tried to stop run to install new TRD alignment
constants and found that run control window is
not responding. Scaler screen showed spill 6,
DAQ monitoring was in spill 71 at 05:16, triggers
running, memory and tape drive lights showing some
activity, and DAQ monitoring showed that data
were going in and out of memory but nothing was
written to tapes. Since we could not stop the run,
we saved online histograms and did
drt_fresh_start.

14643 05:50 Initialized new run with old tapes hoping it would
work and it did. Started new run right after
tevatron quench with 40 min estimate for beam.

06:25 Experts investigating F17 kicker problems. No
estimate for beam.
55 06:58 Beam is back to make the shiftees feel better in
the last two minutes.

=========================================================================
Follow-up from next shift:

Did a muon run and flipped analysis magnet to NEG afterwards.
hcc namelist was loaded fine. No problem!
Constants updated in the DB.

=========================================================================
Follow-up on next shift:
1)Clean out the KTeV refrigerator, before the whole collaboration gets
sick.
2)Look into debugging trigger 16 which is still not working.
3)Order tonner cartridges for the laser printer, Monday day shift.